a school has 42 teachers All the teachers travel to school by car bus or bicycle. 20 of the teachers are male.Five of the male teachers and three female teachers cycle to school. 17 of the teachers travel to school by bus.10 of the female teachers travel to school by car.Complete the two way table to show the numbers of teachers that travel to school by car bus and bicycle.
.........................
Car bus Bicycle Total
Male
Female
Total

Answer by math_tutor2020(3835)


Starting template
CarBusBicycleTotal
Male
Female
Total


The first two pieces of info are:
42 teachers total
20 are male

This means the remaining 42-20 = 22 teachers are female.

We'll have the numbers 20, 22, 42 along the "total" column
CarBusBicycleTotal
Male20
Female22
Total42


Let's move to the next bit of given information:
Five of the male teachers and three female teachers cycle to school

This means we write "5" in the "male" row and "bicycle" column, as well as "3" just directly underneath that "5".
This totals to 5+3 = 8 teachers who travel using a bicycle.
CarBusBicycleTotal
Male520
Female322
Total842


Move onto the next bit of info
17 of the teachers travel to school by bus
We'll write "17" at the bottom of the "bus" column.
CarBusBicycleTotal
Male520
Female322
Total17842


Along the bottom row we have
x+17+8 = 42
where x is the total who drive a car

Solve that equation to get x = 17
There are 17 teachers that use a car
CarBusBicycleTotal
Male520
Female322
Total1717842

Then we're told "10 of the female teachers travel to school by car"
As you can probably guess, we'll have "10" in the "female" row and "car" column.
CarBusBicycleTotal
Male520
Female10322
Total1717842

I'll let the student fill in the remaining empty boxes.